OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -10 - 2 + 6 + 14 + ... + 110

OpenStudy (anonymous):

answer choices : \[\sum_{n=0}^{15}-80n\] \[\sum_{n=0}^{15}-10+8n\] \[\sum_{n=0}^{\infty}-10+8n \] \[\sum_{n=0}^{\infty}-80n\]

OpenStudy (anonymous):

its option B

OpenStudy (anonymous):

i thought so thanks a ton!

OpenStudy (anonymous):

\[\sum_{k=-1}^{14}8k-2\]

OpenStudy (anonymous):

also second answer is true